# Christoforos Kanellakis and George Nikolakopoulos

### v1

@@ I Introduction
Evaluating a traversable route over an actionable 3D scene graph is generally approached in a hierarchical manner, i.e. parsing through corresponding graph representation of abstract layers in a cascading top-down manner. The global semantic-metric path is then shared with a conventional volumetric-based planner to generate a final collision-free path for the robot to track. As scene graphs continuously evolve to support multiple task domain operations such as —such as path planning [5, 7, 6], object search [8], reasoning [9, 10]- their intrinsic composition grows increasingly complex. Intra-layer edges no longer represent only the existence of traversable paths but also encode semantic relationships and affordance mappings between nodes. This results in scene graphs becoming increasingly dense with extraneous information, in terms of computing a path, as the robot explores an environment. Figure. 1 demonstrates an example of multi-domain 3D scene graph, a modified Layered Semantic Graphs (LSG) [11] used in this work to address the task of path-planning. Figure. 1(B-D) highlights the various types of sub-graphs that could be potentially extracted from Figure. 1(A) based on the relevant task.
@@ II-A Contributions
Based on the related state-of-art works in the direction of path planning on 3D scene graphs, we present the technical contributions comprising the proposed SPADE framework in this section.
@@ III SPADE Planner Architecture
Figure. 2 presents a high-level overview of the proposed path planning architecture over 3D scene graphs. As an input to the architecture, SPADE considers the navigational task shared either by an operator as a semantic query, i.e. ``````` ` Inspect open-window in Level-2 of Building-1"""" or directly by the onboard autonomy, i.e. ``````` `GoTo Target-car-3"""", during mission planning. We denote $\textbf{X}_{odom}\in\mathbb{R}^{4}|\textbf{X}_{odom}=[x,y,z,\psi]$ as the localization estimate of the robot provided by the onboard Simultaneous Localization and Mapping (SLAM) module. Furthermore, we define $\mathbb{G}=(V,E)$ as the scene graph entity composed of nodes $v\in V$ and edges $e\in E$ across the abstraction layers $l\in L$. We derive the scene graph representation required for evaluating the proposed architecture from [11]. In [11], the authors define and construct a nested actionable 3D Layered Semantic Graphs (3DLSG) during inspection. The readers are directed to [11] for a detailed understanding of 3DLSG and its construction. Briefly, the 3DLSG in consideration is composed of four main layers, mainly:
@@ VI Conclusions
This work presents the SPADE architecture: a tightly coupled semantic-geometric path planner designed to leverage the ease of initial planning over abstract representations while integrating occupancy based local planner grounded in the volumetric representation of the environment for collision-free navigation. By bifurcating the planning problem depending on the constituent layers, SPADE prioritizes local scene navigation, in the aspect of abstract representation as well as geometric scene model. Furthermore, we demonstrate the impact subsampling multi-domain scene graphs can have on path planning performance with a demonstrated $\approx 40\%$ faster computation of path by a graph-based planner compared to planning on a non-subsampled graph. Additionally, we highlight the adaptability of SPADE to address navigation in dynamic environment conditions that could introduce uncertainty to the traversability condition on previously built 3DLSG. As part of future works, we highlight the integration of SPADE with other defined tasks such as scene reasoning and object search, underscoring the impact of domain-aware subsampling of graphs.

### v2

@@ I Introduction
Evaluating a traversable route over an actionable 3D scene graph is generally approached in a hierarchical manner, i.e. parsing through corresponding graph representation of abstract layers in a cascading top-down manner. The global semantic-metric path is then shared with a conventional volumetric-based planner to generate a final collision-free path for the robot to track. As scene graphs continuously evolve to support multiple task domain operations such as path planning [5, 7, 6], object search [8], reasoning [9, 10], their intrinsic composition grows increasingly complex. Intra-layer edges no longer represent only the existence of traversable paths but also encode semantic relationships and affordance mappings between nodes. This results in scene graphs becoming increasingly dense with extraneous information, in terms of computing a path, as the robot explores an environment. Figure. 1 demonstrates an example of multi-domain 3D scene graph, a modified Layered Semantic Graphs (LSG) [11] used in this work to address the task of path-planning. Figure. 1(B-D) highlights the various types of sub-graphs that could be potentially extracted from Figure. 1(A) based on the relevant task.
@@ III SPADE Planner Architecture
Figure. 2 presents a high-level overview of the proposed path planning architecture over 3D scene graphs. As an input to the architecture, SPADE considers the navigational task shared either by an operator as a semantic query, i.e. ``‘ ‘ Inspect open-window in Level-2 of Building-1" or directly by the onboard autonomy, i.e. ``‘ ‘GoTo Target-car-3", during mission planning. We denote $\textbf{X}_{odom}\in\mathbb{R}^{4}|\textbf{X}_{odom}=[x,y,z,\psi]$ as the localization estimate of the robot provided by the onboard Simultaneous Localization and Mapping (SLAM) module. Furthermore, we define $\mathbb{G}=(V,E)$ as the scene graph entity composed of nodes $v\in V$ and edges $e\in E$ across the abstraction layers $l\in L$. We derive the scene graph representation required for evaluating the proposed architecture from [11]. In [11], the authors define and construct a nested actionable 3D Layered Semantic Graphs (3DLSG) during inspection. The readers are directed to [11] for a detailed understanding of 3DLSG and its construction. Briefly, the 3DLSG in consideration is composed of four main layers, mainly:
@@ VI Conclusions
[…]